### v2

# Pressure-dependent Electronic Superlattice in the Kagome-Superconductor CsV3Sb5

Abstract
We present a high-resolution single crystal x-ray diffraction study of kagome-superconductor CsV3Sb5, exploring its response to variations in pressure and temperature. We discover that at low temperatures, the structural modulations of the electronic superlattice, commonly associated with charge-density-wave order, undergo a transformation around $p\sim$ 0.7 GPa from the familiar $2\times 2$ pattern to a long-range-ordered modulation at wavevector $q=(0,3/8,1/2)$. Our observations align with inferred changes in the CDW pattern from prior transport and nuclear-magnetic-resonance studies, providing new insights into these transitions. Interestingly, the pressure-induced variations in the electronic superlattice correlate with two peaks in the superconducting transition temperature as pressure changes, hinting that fluctuations within the electronic superlattice could be key to stabilizing superconductivity. However, our findings contrast with the minimal pressure dependency anticipated by ab initio calculations of the electronic structure. They also challenge prevailing scenarios based on a Peierls-like nesting mechanism involving van Hove singularities.
Understanding the correlation between unconventional superconductivity (SC) and spatial electronic modulations, encompassing magnetic, nematic, or charge-density-wave (CDW) orders, represents one of the key challenges in contemporary condensed matter research [1, 2, 3]. Consequently, extensive endeavors are underway to unravel the underlying physics in quantum materials like cuprates [4], nickelates [5, 6], iron or nickel pnictides [7, 8, 9], transition metal dichalcogenides [10] or heavy fermion compounds [11], where the relation between SC and CDW phenomena can be examined in different settings.
In this framework, a new and extremely intriguing class of materials has recently been discovered: the layered kagome compounds AV3Sb5 (A=K, Cs, Rb) [12]. Characterized by an electronic structure featuring flat bands, van Hove singularities (vHs), Dirac cones, and non-trivial band topology, these materials offer a unique platform for exploring novel electronic states of matter with intertwined orders [3]. Particularly noteworthy is the detection of an electronic superlattice below 94 K, [12] which is commonly associated with a CDW, and that is succeeded by superconductivity below $T_{c}\simeq 2.7$ K in CsV3Sb5– a revelation that has sparked considerable excitement. This excitement has been further substantiated by recent accounts of the absence of Kohn anomalies in the phonon dispersion associated with the formation of the CDW [13, 14], chirality [15], time-reversal symmetry breaking or electronic nematicity [16] of the electronic superlattice. Consequently, while we refer to the electronic modulation in CsV3Sb5 as a CDW in the following, we recognize that it is often rightly considered unconventional and that the underlying electronic order may be substantially more complex. As a matter of fact, numerous important aspects of this CDW and of its relationship with SC in CsV3Sb5 remain elusive and subject to intense debate [16, 17, 18, 19, 20]. The origin of the CDW and the nature of the interplay between electronic and lattice degrees of freedom in these materials therefore requires further elucidation [13, 21, 14, 22].
Figure 1: Temperature and pressure dependence of the superlattice modulation observed for sample $K$ (top) and $D$ (bottom). $hl$- and the $hk$-cuts through XRD intensity distributions in the reciprocal space are shown. At 40 K and $p=0.23$ GPa, the two samples both display a $2\times 2$ modulation along $ab$ but a different superstructure along $c$. Increasing the pressure at 40 K above 0.7 GPa, here shown for 1.14 GPa, induces a new 3/8-CDW-order, which is the same for both samples. The green dashed lines in panel (a) indicate the $2\times 2$x4 peaks, in panel (b) the $2\times 2$x2 peaks, and for (c) and (d) the 3/8-peaks.
To address such issues, pressure tuning has proven to be a valuable tool in quantum materials research [23, 24, 25], providing a reversible method to adjust the balance between competing energy scales and enabling the investigation of complex phase diagrams in these materials without introducing chemical disorder. This approach is without doubt most relevant in the case of kagome superconductors, since both SC and CDW transition temperatures have been reported to strongly depend on hydrostatic pressure, displaying intriguing behavior already for relatively modest pressures [26, 27, 17]. This is demonstrated across a diverse array of measurements encompassing electrical- and magnetotransport, magnetic susceptibility, nuclear magnetic and quandrupole resonances (NMR and NQR), x-ray diffraction (XRD), and muon spin resonance [26, 27, 17, 28, 29] which have revealed that both the superconducting transition temperature ($T_{c}$) and the upper critical field $H_{c2}$ as a function of pressure $p$ exhibit a distinctive double-dome feature below $p<2.0$ GPa , characterized by a first maximum of $\sim$ 7 K at $p_{1}\simeq 0.7$ GPa, succeeded by a second maximum at $p_{2}\simeq 2$ GPa with $T_{c}\simeq$ 8 K [17, 27, 30, 31]. The CDW transition temperature appears monotonically suppressed with pressure, but from a marked reduction of the magnetoresistance [26] and the appearance of new NMR lines above $p_{1}$, changes in the electronic modulation pattern have been indirectly inferred.
Nonetheless, a thorough examination of the pressure dependence of CDW superstructures across various temperatures – essential for clarifying the microscopic physics at play – has yet to be undertaken. Here we close this gap, by reporting a detailed pressure- and temperature-dependent single crystal x-ray diffraction (XRD) study of the crystal lattice and its superstructures in CsV3Sb5. We report the gradual suppression of the familiar $2\times 2$ CDW modulation with pressure, alongside the appearance of a new type of modulation with a completely different pattern and periodicity above $p_{1}$. The two orders coexist for a while before the disappearance of the former. At 25 K, no CDW survives above $\sim$ 1.7 GPa.
The clear change of the CDW pattern we report in the investigated pressure range, however, contrasts with the modest impact of pressure on the electronic structure as computed – using the experimentally determined pressure-dependent lattice parameters – by first-principles calculations. In particular, the van Hove singularities below the Fermi energy show minimal pressure dependence in the investigated range. This discrepancy challenges weak coupling scenarios that attribute the formation of electronic superlattices in kagome metals AV3Sb5 (A=K, Cs, Rb) to a Peierls-like nesting of van Hove singularities [32, 33].
Figure 2: Pressure dependence of the CDW order in CsV3Sb5. The contour plot shows the XRD-intensity along the $A-L$ reciprocal direction of the hexagonal cell (obtained from the averaging of the intensity measured at 25 K along the six reciprocal $A-L$ directions from the (3, 0, 0.5) A point). The intensity at the (3.5, 0, 0.5) point (L-point of the hexagonal cell) below 0.7 GPa signals the $2\times 2\times 2$ order, which transforms into the 3/8 modulation with increasing pressure revealed by the intensity at $q=3/8$ and $q=5/8$ above 0.7 GPa.
We investigated single crystals from two distinct batches, referred to as batch $D$ and batch $K$ in the following, synthesized and characterized, as described in the supplementary information [34]. High-resolution XRD measurements as a function of $T$ and $p$ were performed at the ID15B beamline of the ESRF in Grenoble, France, using a monochromatic x-ray beam with an energy of 30.17 keV, focused down to 2x4 $\mu$m2 at the sample position. The diffraction patterns were measured with an EIGER2 X 9M CdTe flat panel detector. Diamond anvil cells (DACs) with helium as a pressure transmitting medium were used to apply pressures of up to 27 GPa under close to hydrostatic conditions. In each DAC, one sample of batch $D$ and one of batch $K$ were loaded into the sample space, so that both samples were in exactly the same experimental condition during our pressure and temperature dependent XRD measurements. Further details are given in the supplementary information [34].
At room temperature and ambient pressure, both samples exhibit a hexagonal $P6/mmm$ structure, in full agreement with previously published results [35]. In the following, reflections will be always indexed referring to this structure, even though there are structural transitions as a function of $p$ and $T$ [34]. In agreement with previous reports [12], both samples develop a $2\times 2$ superstructure within the $ab$-planes when cooled below 94 K at ambient or small pressure. This CDW transition is signalled by additional satellite reflections, as demonstrated in panels (a) and (c) of Fig. 1. Interestingly, as can also be observed in these panels, the structural modulation of the two set of samples differs along the $c$-direction. At 40 K and $p=0.23$ GPa, despite special care taken to ensure they are precisely in the same experimental condition and sharing the same $pT$-history, sample $K$ exhibits a $2\times 2\times 4$ modulation, whereas sample $D$ displays a $2\times 2\times 2$ order. This observation supports earlier conclusions that the orders along $c$ are metastable [36], while the $2\times 2$ modulation within the $ab$-layers is a defining feature of the electronic ordering instability of CsV3Sb5 at ambient pressure [37]. In both cases, the superstructure reflections are best indexed as single-Q modulations of an orthorhombic $Cmmm$ phase [34], of symmetry lower than the original $P6/mmm$ in agreement with findings of Ref. 38.
The major experimental finding of our study is represented by the data in the right panels (b) and (d) of Fig. 1: Irrespective of the periodicity of the modulation along the $c$-axis, an increase of the pressure $p$ to 0.7 GPa at constant T=40 K induces in both samples a transition into a new phase (see also Supplementary Fig. 8 [34]) characterized by the same commensurate modulation vector $q=(0,\frac{3}{8},\frac{1}{2})$. We will refer to this as the ’3/8’ phase in the following. This observation aligns very well with recent findings from ${}^{51}V$ NMR [17] and ${}^{121/123}Sb$ NQR [27], which reported in this very pressure range the appearance of new spectral features that are distinct from those of the $2\times 2$ modulation and have been interpreted as evidence of a new type of CDW order. Although these local probes cannot directly determine the ordering wave vector, proposals for an incommensurate structure [27] or a unidirectional order with a period of $4\times a$ [17] have been suggested. While our XRD data confirms a pressure-driven change in the electronic order, we unambiguously reveal here that the $p$-induced ordering is commensurate and different from the suggested $4\times a$.
With a further increase in $p$ at constant $T$, the intensities of the superlattice peaks of the 3/8-phase gradually diminish and vanish above 1.7 GPa, again in good agreement with the nuclear resonance data [17, 27]. According to our structural refinement, the 3/8 phase is accompanied with a small monoclinic distortion [34].
To further elucidate the evolution of the CDW order with respect $p$, we show in Fig. 2 the pressure dependence of the XRD-intensity along the high symmetry $A-L$ line of the hexagonal cell (starting from the reciprocal A point at (3, 0, 0.5)) as a function of $p$ at constant $T=25$ K. With $p$ increasing towards 0.7 GPa, the satellite reflections of the $2\times 2$ in-plane order are gradually suppressed. As can be observed in Fig. 2, the intensity of this peak weakens significantly around 0.7 GPa, but remains detectable up to 0.9 GPa. At the same time, the new 3/8 modulation emerges around 0.7 GPa, i.e., the $2\times 2$ and the 3/8 phases coexist between 0.7 and 0.9 GPa. Beyond 0.9 GPa, the $2\times 2$ disappears completely, and by further increasing $p$ also the scattering due to the 3/8 modulation is progressively reduced until it falls below our detection limit at 1.7 GPa. The data shown in Fig. 2 further reveals that, within the error of our experiment the positions of the satellite peaks corresponding to the $2\times 2$ and the 3/8 order do not change with varying $p$ and that, despite the strong suppression of the intensity, the satellite reflections always remain as sharp as the Bragg reflections and are essentially resolution limited [34].This is true both for in- and out-of-plane directions.
Figure 3: Electronic phase diagram of CsV3Sb5 in the low-pressure and low-temperature region. The orthorhombic $2\times 2\times 2$ and $2\times 2\times 4$ phases, summarized as $2\times 2$, exist in the green region below $\approx 0.7$ GPa. The new monoclinic phase with the 3/8-order is indicated in blue. Both phases coexist in a narrow pressure range as indicated by the striped area. The slope at low pressure ($dT_{CDW}/dp\sim-75K/GPa$) follows that dertermined from thermodynamic measurements [19]. The red dashed line sketches the evolution of the superconducting $T_{c}$ as a function of pressure (after Ref. 17). Inset: shows the hexagonal vs. orthorhombic lattice in the kagome plane.
The results of our XRD studies below 100 K and up to 2.5 GPa are summarized in the $pT$-phase diagram presented in Fig. 3, where the stability regions of the $2\times 2$ and the 3/8 orders are indicated by the green and blue areas, respectively. Comparing the pressure dependence of the electronic order to the pressure dependence of the superconducting $T_{c}$ [39, 40] reveals an interesting correlation: The pressure $p_{1}\simeq 0.7$ GPa, where $T_{c}(p)$ shows its first maximum, aligns perfectly to the critical pressure where the $2\times 2$ order is suppressed and the 3/8 order appears. Furthermore, the pressure $p_{2}\simeq 2$ GPa of the second maximum of $T_{c}(p)$ coincides well with that at which the 3/8 phase disappears.
The observation that every time an electronic order is suppressed with increasing $p$, the superconducting transition temperature $T_{c}(p)$ exhibits a maximum is remarkable. One obvious explanation to be considered is, of course, the competition between CDW and SC. Direct experimental evidence for this competition, for example in the form of a suppression of the CDW in the superconducting state is still lacking. Nevertheless, the observation of a peak in $T_{c}(p)$ near the critical point of a distinct electronically ordered phase (here the $2\times 2$ or the 3/8 phase) suggests that the corresponding fluctuations may contribute to supporting superconductivity. This conjecture certainly warrants careful examination in future investigations.
It is important to note that the consistently resolution-limited satellite reflections (cf. Fig. 2) do not necessarily imply the absence of fluctuations. The resolution-limited width of these reflections simply indicates that the coherent contribution to a given satellite reflection originates from a volume of about $10^{4}$ unit cells [34]. However, it remains indeterminate whether this volume constitutes a single contiguous entity or a series of smaller, distinct volumes maintaining a fixed phase relation (see Ref. [41]). In essence, despite the sharpness of the satellite peaks, regions experiencing fluctuations may still exist. These fluctuations would manifest as an exceedingly weak diffuse scattering signal, beyond the detection capabilities of the current experimental setup.
The comparison of the present $p$-dependent XRD results to those of a recent doping dependent study [42] is also very interesting: In the doping dependent study, it was observed that replacing Sb with Sn destabilizes the $2\times 2$ superlattice as well, leading to the emergence of a new superlattice at low temperatures with a modulation wave vector similar to that of the 3/8 phase observed here. However, there are notable distinctions between the doping and $p$-induced orders: The doping-induced phase exhibits short and highly anisotropic correlation lengths within the $ab$-layers that have been interpreted as signs of electronic nematicity. In contrast, the the resolution-limited superlattice modulations observed in pristine CsV3Sb5, including the pressure-induced 3/8 phase, did not reveal any detectable in-plane anisotropy.
In principle, Sn-substitution and $p$ do not represent equivalent modifications of CsV3Sb5. On a qualitative level, however, they both affect the $c/a$ ratio, which has been found to play a critical role for controlling the electronic properties [19, 22]. Due to the anisotropic layered structure of CsV3Sb5, hydrostatic pressure has indeed a pronounced effect on the c/a ratio, which can be determined precisely by our measurements. We find that a pressure as modest as 1 GPa is sufficient to reduce c/a by about 3% (cf. Supplementary Fig. 4 [34]).
[…]
The drastic pressure dependence of the electronic superlattice reported here therefore implies that the nesting of vHs cannot be the primary stabilization mechanism in CsV3Sb5. If this was the case, the $2\times 2$ order would be expected to remain entirely stable below 1 GPa, at odds with our experimental observation. This conclusion is further supported by a detailed analysis of the Lindhard function from DFT, which shows that the electronic system alone does therefore not favor a specific ordering wave vector [45]. Even though the new 3/8 phase has not been predicted by any first principles calculations to the best of our knowledge, it would be essential to check within DFT whether it could become more stable than a $2\times 2$ CDW under pressure. Strong coupling scenarios involving strongly momentum-dependent electron-phonon coupling [46, 45, 47] or lattice anharmonicity effects [48] have been considered as a driving force behind the stabilization of a specific charge order instabilities in these materials. Along these lines, further advanced first-principles calculations of the temperature and pressure dependence of the phonon spectrum and of the electron-phonon coupling might be required to clarify the mechanism behind the dramatic pressure-induced change of modulation reported here.
In conclusion, the present study yields two major findings namely the correlation between the critical pressures of the electronic modulations and superconductivity in CsV3Sb5 and a transition from the $2\times 2$ to a long-range charge ordered 3/8 phase at pressures below 1 GPa. Furthermore, our findings reveal pressure-induced structural modulations that appear to coincide with the trend in the superconductivity manifested in other experiments. These observations strongly substantiate the notion that critical fluctuations and complex electron-lattice interactions, extending well beyond conventional scenarios and models, might play a prominent role in shaping the phase diagram of AV3Sb5.
Acknowledgements
We acknowledge fruitful discussions with R. Fernandes, R. Heid, M. Merz and S. M. Souliou. This research received support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), under projects 422213477 (TRR 288 project B03) and 247310070 (CRC 1143 project C06 and B05). T.R., F.S., C.F., and J.G. express gratitude to the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter (ct.qmat EXC-2147, Project No. 390858490) for financial backing. Additionally, we extend our appreciation to the European Synchrotron Radiation Facility (ESRF) for providing access to their synchrotron radiation facilities.